OpenStudy (anonymous):

someone please help me. i have been trying to figure this out. 1.Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

(0.50+0.2)x = 500 (im assuming you mean 40% of EACH flower's price so i did 0.50x0.4=0.2)

OpenStudy (anonymous):

well that part i do get. you have to do 40% of .50. my teacher said part of it i got right. i know they would have to sell 2500 flowers in order to make 500 dollars. they only part i dont get is how to set the equasion up.

OpenStudy (anonymous):

okay then how did you calculate 2500?

OpenStudy (anonymous):

because the price of each flower equals 0.20 and i multiplied that by 500

OpenStudy (anonymous):

thats 100

OpenStudy (anonymous):

sorry i said that wrong

OpenStudy (anonymous):

i cant remember how i got 2500 but my teacher said it was right. each flower costs 0.20

OpenStudy (anonymous):

strange...if i understood the question, the equation is: (0.5+0.2)x - 0.5x = 500 the terms in the bracket represent the sale price of each flower and the -0.5x is the cost to buy each flower

OpenStudy (anonymous):

yes, but if you do 0.20 and multiply it by 2500 then you get 500

OpenStudy (anonymous):

that can only mean that the equation has to be 500/0.2 to give 2500 and i dont see a way that would work...either you misunderstood your teacher or i am misunderstanding the question lol. sorry, wish i could have helped

OpenStudy (anonymous):

thats ok thanks for trying.